Solve


(D^(2)+DD'-6D'^(2))z=y cosx


1
Expert's answer
2022-01-09T16:55:09-0500

Answer:


Given:



"(D ^\n2\n +DD \n^\u2032\n \u22126D \n^{\u20322}\n )z=ycosx"



Now, the Auxiliary equation will be:


"m \n^2\n +m\u22126=0\n\u27f9(m\u22122)(m+3)=0\n\u27f9m=2,\u22123"


Hence,

"C.F=\u03d5_ \n1\n\u200b\n (y+2x)+\u03d5 \n_2\n\u200b\n (y\u22123x)"

Now,


"P.I=\frac{1}{(D ^
2
 +DD 
^′
 −6D 
^{′2}
 )}ycosx"



"\implies P.I=\frac{1}{(D+3D^{'})(D-2D^{'})}ycosx=\frac{1}{D+3D^{'}}\int_{D-2D^{'}}ycosxdx=\frac{1}{D+3D^{'}}[(c_1-2x)sinx-2cosx]=\frac{1}{D+3D^{'}}(ysinx-2cosx)=\int_{D+3D^{'}}(ysinxdx)-\int_{D+3D^{'}}2cosxdx=\int _{D+3D^{'}}(c_2+3x)sinxdx=\int _{D+3D^{'}}2cosxdx"

"(y=c_2+3x)"

"\implies P.I=-ycosx+\int cosxdx = -ycosx +sinx"

Therefore the complete solution is


"C.F+P.I= \phi_1(y_+2x)+\phi_2(y-3x)+-ycosx+sinx"